OpenStudy (anonymous):

A grocery store is selling 12 bananas for $2.40. At this rate, what is the cost of 15 bananas? A. $2.52 B. $2.55 C. $3.00 D. $36.00

OpenStudy (anonymous):

b

OpenStudy (black_ninja123):

Do 2.40/12 first

OpenStudy (texaschic101):

you could also do it by setting it up as a proportion

OpenStudy (black_ninja123):

You will get 0.2 and than multiply 0.2*15=3 So your answer is C

OpenStudy (texaschic101):

12/2.40 = 15/x cross multiply 12x = (2.40)(15) 12x = 36 x = 36/12 x = 3 I agree...answer is C

OpenStudy (anonymous):

I AGREE TOO.